OpenStudy (anonymous):

For the unbalanced equation NH3 + O2 → NO + H2O If at the beginning of the reaction you used 4.30 g NH3 and 15.76 g O2, what is the theoretical yield (in grams) of H2O (you will need to find the limiting reagent and then the theoretical yield of H2O from this reagent)?

OpenStudy (photon336):

You need to balance the equation first before doing anything else

OpenStudy (anonymous):

what Do I have to do after I Have balanced the equation??

OpenStudy (photon336):

Calculate the molar mass of NH3 O2 and H2O Could you show me what you have so far?

OpenStudy (anonymous):

NH3 - 4.30g X 4 mol/17.03g = 1.009982384 mols

OpenStudy (anonymous):

02 - 15.76g* 5 mol/32g = 2.4625 h20- 6 mol * 18.02/1 mol = 108.12 g